# EO-1: Interleaved Vision-Text-Action Pretraining for General Robot Control

### v4

@@ Abstract
The human ability to seamlessly perform multimodal reasoning and physical interaction in the open world is a core goal for general-purpose embodied intelligent systems. Recent vision-language-action (VLA) models, which are co-trained on large-scale robot and visual-text data, have demonstrated notable progress in general robot control. However, they still fail to achieve human-level flexibility in interleaved reasoning and interaction. In this work, introduce EO-Robotics, consists of EO -1 model and EO -Data1.5M dataset. EO -1 is a unified embodied foundation model that achieves superior performance in multimodal embodied reasoning and robot control through interleaved vision-text-action pre-training. The development of EO -1 is based on two key pillars: (i) a unified architecture that processes multimodal inputs indiscriminately (image, text, video, and action), and (ii) a massive, high-quality multimodal embodied reasoning dataset, EO -Data1.5M, which contains over 1.5 million samples with emphasis on interleaved vision-text-action comprehension. EO -1 is trained through synergies between auto-regressive decoding and flow matching denoising on EO -Data1.5M, enabling seamless robot action generation and multimodal embodied reasoning. Extensive experiments demonstrate the effectiveness of interleaved vision-text-action learning for open-world understanding and generalization, validated through a variety of long-horizon, dexterous manipulation tasks across multiple embodiments. This paper details the architecture of EO -1, the data construction strategy of EO -Data1.5M, and the training methodology, offering valuable insights for developing advanced embodied foundation models.
@@ 1 Introduction
Generalist robot policies with open-world capabilities is essential for deploying autonomous robots in real scenarios, where they need to tackle diverse tasks, follow various instructions, adapt to different situations, and even handle unforeseen events. In the realm of digit domains, the paradigm of learning from large-scale multimodal datasets has demonstrated broad proficiency and generalization across various interactive and assistive applications. Extending this idea to the physical world, researchers are now training large vision-language-action models on extensive robotic datasets to realize generalist robots with similar versatility. Although recent efforts have demonstrated impressive progress in achieving dexterous manipulation and long-horizon physical control within constrained environments, open world generalization remains the central challenge for developing general-purpose autonomous AI agents in the physical world (Black et al., 2025, Gemini-Robotics, 2025). To address this challenge, robots must be capable of acquiring comprehensive world knowledge, engaging in human-level reasoning, and executing dexterous actions.
@@
Early generalist robot policies (Kim et al., 2024, Black et al., 2024, Pertsch et al., 2025) primarily extend vision–language models (VLMs) into vision-language-action (VLA) models with domain-specific robotic data, either through auto-regressive decoding of discrete action tokens or by incorporating additional continuous flow matching modules. However, because these VLA models are trained exclusively on robotic datasets, they are restricted to narrow task domains and specific environments. As a result, they suffer from diminished general semantic knowledge inherited from VLMs and exhibit limited instruction-following capabilities.
@@
Recently, several studies (Black et al., 2025, Driess et al., 2025, Lin et al., 2025) have explored co-training VLA models with both web data and robotic data, showing promising generality when interacting with new objects and unseen backgrounds. Nevertheless, existing approaches remain predominantly generating robot action at the end of the VLA output sequence, overlooking the rich temporal dynamics and causal dependencies among vision, language, and action modalities inherent in open-world embodied interactions. Humans, in contrast, exhibit a flexible and interleaved synergy between multimodal embodied reasoning and physical action, enabling highly generalizable and dexterous manipulation in the open world, i.e., reasoning guides action and action results inform subsequent reasoning. This raises a fundamental research question: How can we design an effective training paradigm for generalist robot policies that support flexible and mutual-informed reasoning-acting integration?
@@
Inspired by the impressive results of advanced multimodal understanding and generation systems (Deng et al., 2025, Ma et al., 2025b, Xie et al., 2025, NextStep-1, 2025), which demonstrate the superiority in interleaved multimodal pretraining, we propose a unified embodied model to enable flexible and powerful multimodal embodied reasoning and action generation through interleaved embodied pretraining. Achieving this requires not only an effective and unified architecture that excels in mix-modality generation, but also carefully structured multimodal embodied data that jointly integrate texts, images, videos, and robotic actions.
@@
Regarding model architecture, we adopt a single unified decoder-only transformer that integrates discrete auto-regressive decoding with continuous flow matching denoising. The unified model is built on a pre-trained VLM, thereby inheriting broad visual–language knowledge, and its shared parameters are further optimized with modality-specific objectives: next-token prediction for text and flow matching for robotic actions. The model applies causal attention cross the entire interleaved vision-text-action sequence to capture the sequential dependencies between reasoning and acting. In addition, two MLPs are added to encode and decode continuous robotic actions, complementing the original text and visual tokenizers. Unlike prior VLA models (Black et al., 2024, 2025, Driess et al., 2025) that introduce extra action-specific modules to learn action generation, our design enables easier alignment between vision/language and action modalities through circumventing train new action-specific parameters from scratch, thereby achieving more effective cross-modal knowledge transfer in generalist robot policies.
@@ 2.1 Problem Formulation
where $V_{\pi_{\theta}}$ represents the parameters of $\pi_{\theta}$ to predict the vector field for robot action generation. The actions are generated by integrating the predicted vector field from $\tau=0$ to $\tau=1$, starting with random noise ${z}_{t}^{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and $\delta$ is the integration step size. The shared parameters in the unified transformer enable the seamless transfer of semantic knowledge from vision-language understanding to action generation, leading to more general reasoning capabilities and control capabilities in embodied scenarios (Deng et al., 2025, Black et al., 2025, Driess et al., 2025).
@@ 2.2 Model and Training
The proposed unified model architecture processes interleaved multimodal inputs (i.e., text, images, videos, and actions) through a shared transformer backbone that generates both discrete token sequences and continuous action tokens. The model employs a text tokenizer and visual encoder to convert text and image patches into input tokens, while the robot state $\mathbf{q}_{t}$ is linearly projected into the same transformer embedding space, i.e., $x_{i}^{w},x_{i}^{I},x_{i}^{q}\in\mathbb{R}^{d}$. Note that the text tokenizer and visual encoder are inherited from pre-trained VLM and the state projector is random initialized. For flow matching action denoising, the input “noisy action” $\mathbf{a}_{t}^{\tau}$ is obtained by $\mathbf{a}_{t}^{\tau}=\tau\mathbf{a}_{t}+(1-\tau)z_{\tau}$, where $z_{\tau}\sim\mathcal{N}(0,I)$ is random noise. Then, another noisy action linear projector is utilized to embed noisy action and flow matching timestep $\tau$ into noisy action token $x_{i}^{a}\in\mathbb{R}^{d}$. The unified transformer backbone, initialized from Qwen 2.5 VL (Bai et al., 2025), processes this interleaved multimodal input sequence with shared parameters and generates output sequence through two separate heads: a language head for text token decoding and a flow head for continuous action denoising. Note that we do not introduce extra action-specific parameters to separately model action denoising as prior VLA models, enabling seamless integration of discrete multimodal embodied reasoning and continuous robot control.
@@
where $\mathbf{a}_{t}^{\tau}=\tau\mathbf{a}_{t}+(1-\tau)z_{0}$ is the input noisy action. $x_{<a}$ represents the all preceding multimodal tokens. As in (Black et al., 2024), we sample the flow matching timestep $\tau$ from a beta distribution that emphasizes lower (noisier) timesteps. With the interleaved multimodal token sequence, we apply next token prediction $\mathcal{L}_{ar}(\theta)$ to the language head and flow matching $\mathcal{L}_{fm}(\theta)$ to the flow head for predicting denoising vector field. The unified model is trained end-to-end by optimizing sum of these two objectives: $\mathcal{L}=\mathcal{L}_{ar}(\theta)+\mathcal{L}_{fm}(\theta)$.
@@ 3 Dataset and Benchmark
LLaVA-1.5 (Liu et al., 2024a), LLaVA-Video-178K (Zhang et al., 2024b),
PixMo-Points (Deitke et al., 2024a), RefCOCO (Kazemzadeh et al., 2014a),
RoboVQA (Sermanet et al., 2024)
@@
AgiBotWorld (Bu et al., 2025), Open X-Embodiment (OpenX-Embodiment et al., 2024),
RoboMIND (Wu et al., 2024), SO100-Community (Shukor et al., 2025),
IPEC-Franka (Qu et al., 2025)
@@ 3.1 Data Source
Web Multimodal Data. The web text-image paired data is essential for multimodal understanding. We curated a web multimodal dataset comprising 5.7 Million samples with 7.1 Billion tokens, integrating six public sources across three task domains: i) Visual instruction following. We incorporate the LLaVA Visual Instruct Series (Liu et al., 2023b, Sermanet et al., 2024) datasets, which contain GPT-generated multimodal instructions and robotics-oriented visual question answering examples. These sources cover diverse interaction patterns, including image-grounded conversation, multi-step reasoning, and long-horizon task planning. ii) Video understanding and reasoning. We leverage the LLaVA-Video (Zhang et al., 2024a) and RoboVQA (Sermanet et al., 2024), which support temporal multimodal comprehension through detailed video captioning, open-ended QA, and multiple-choice question answering. iii) Referring expression and spatial grounding. To support fine-grained object understanding, we include RefCOCO (subset) (Kazemzadeh et al., 2014b) for natural language referring expression comprehension and PixMo-Points (Deitke et al., 2024b) for spatial localization via human-annotated coordinate grounding.
Robot Control Data. Robot control data provides critical supervision for policy learning across embodiment generalization and fine-grained manipulation. We aggregate a large-scale real robot control dataset comprising 1.2 Million episodes with 0.13 Trillion tokens from five public sources, forming one of the most comprehensive collections for robot learning to date. The dataset includes AgiBot-World (Bu et al., 2025), featuring dual-arm humanoid demonstrations across diverse bimanual manipulation tasks with rich visuomotor sequences; Open X-Embodiment (OpenX-Embodiment et al., 2024), which spans 22 robot types and 527 skills to support cross-platform generalization; and SO100-Community (Shukor et al., 2025), a collection of long-horizon community demonstrations using SO100 platform. It also incorporates RoboMind (Wu et al., 2024), covering a wide range of tasks across multiple embodiments (e.g., Panda, UR5e, AgileX, humanoids) with diverse object classes.
@@ 4.1 Implementation Details
The EO -1 model is trained on a large-scale corpus that integrates 1.2M real-robot demonstrations from AgiBotWorld (Bu et al., 2025), Open X-Embodiment (OpenX-Embodiment et al., 2024), RoboMIND (Wu et al., 2024), and SO100-Community (Shukor et al., 2025), together with 5.7M web multimodal samples and 1.5M interleaved embodied data pairs, yielding a total of 135B tokens. To optimize on this scale, we train for five epochs with Flash-Attention variable-length packing (average sequence length of 16,384) and a batch size of 1. The backbone learning rates are set to $5\times 10^{-5}$ for both the language model and multimodal projector, $1\times 10^{-5}$ for the vision encoder, and a fixed balance factor of $1.0$ is applied between language regression and action flow matching losses. Additional settings include an action denoising chunk size of $16$, resolution bounds of 50176–100352, and the DeepSpeed ZeRO-1 optimizer. At inference time, the policy is conditioned on multi-view camera observations and sub-task instructions. It predicts a 16-step action chunk through 10 denoising iterations for robot execution. This design achieves a balance between stability and efficiency, while requiring only 6 GB of GPU memory on a single NVIDIA RTX 4090, thus enabling real-time evaluation in both simulation and real-world environments.
@@ 4.2 Open-sourced Benchmark Evaluation Setups
  • RoboVQA (Sermanet et al., 2024): A free-form visual question answering (VQA) benchmark designed to assess high-level reasoning in egocentric embodied scenes. Following prior works, model performance is evaluated using BLEU-4 scores.
  • ERQA (Gemini-Robotics, 2025): A curated benchmark that emphasizes spatial reasoning and grounded world knowledge. The tasks are designed around realistic robotic scenarios requiring inference over object relationships, geometry, and agent-centric spatial cues.
@@
  • SimplerEnv (Li et al., 2024b): A manipulation benchmark featuring WidowX and Google Robots operating in visually diverse environments with variations in lighting, surface textures, and camera viewpoints. It leverages the Bridge Data V2(Walke et al., 2023) for evaluating Real-to-Sim transfer. SimplerEnv includes both short-horizon atomic tasks and visually challenging variants, specifically designed to assess robustness under appearance shifts between real and simulated settings.
  • LIBERO (Liu et al., 2023a): A long-horizon manipulation benchmark consisting of temporally extended, multi-stage tasks across various object categories and interaction types. LIBERO emphasizes planning, memory, and semantic understanding in complex simulated environments.
@@ 4.3 Benchmark Evaluation Results
Embodied Reasoning. To evaluate the embodied reasoning capabilities, we benchmark it on three key tasks: RoboVQA Sermanet et al. (2024), ERQA (Gemini-Robotics, 2025), and EO -Bench. These benchmarks assess a diverse range of reasoning skills, including spatial understanding, physical commonsense, and multistep task planning. We compare EO -1 against three categories of baselines: public VLMs (e.g., Qwen2.5 VL (Bai et al., 2025), InternVL2.5 (Chen et al., 2024b)), private VLMs (e.g., GPT-4o, Gemini 1.5 Flash (Gemini et al., 2023b), Claude 3.5), and co-trained visual-language-action models (e.g., ChatVLA (Zhou et al., 2025), Magma (Yang et al., 2025)), with all models evaluated under consistent embodied conditions.
@@
Gemini 1.5 Flash (Gemini et al., 2023a) 46.0 46.3 46.3 37.8 44.1
@@
Qwen2.5 VL 3B (Bai et al., 2025) 55.9 35.3 20.0 22.6 33.5
Qwen2.5 VL 7B (Bai et al., 2025) 56.9 39.3 26.7 31.5 38.6
InternVL2.5 4B (Chen et al., 2024a) 45.7 36.8 37.0 31.9 37.9
InternVL2.5 8B (Chen et al., 2024a) 43.9 45.2 32.8 38.1 40.0
ChatVLA 2B (Zhou et al., 2025) 33.7 34.3 26.4 21.9 29.1
Magma 8B (Yang et al., 2025) 30.3 29.3 29.4 36.7 31.4
@@
Table 2: Performance comparison with standard VLMs and co-trained visual-language-action models on RoboVQA (Sermanet et al., 2024), ERQA (Gemini-Robotics, 2025), and EO -Bench benchmarks.
@@
OpenVLA (Kim et al., 2024) 84.7 0.9% 5 88.4 0.8% 5 79.2 1.0% 4 53.7 1.3% 5 76.5 0.6% 5
SpatialVLA (Qu et al., 2025) 88.2 0.5% 2 89.9 0.7% 4 78.6 0.6% 5 55.5 1.0% 4 78.1 0.7% 4
OpenVLA-OFT (Kim et al., 2025) 97.6 0.9% 1 98.4 0.8% 1 97.9 1.0% 1 94.5 1.3% 1 97.1 0.6% 1
ThinkAct (Huang et al., 2025) 88.3 –% 3 91.4 –% 3 87.1 –% 3 70.9 –% 3 84.4 –% 3
MolmoAct-7B-D (Lee et al., 2025) 87.0 –% 4 95.4 –% 2 87.6 –% 2 77.2 –% 2 86.6 –% 2
Diffusion Policy (Chi et al., 2023) 78.5 1.1% 6 87.5 0.7% 5 73.5 1.2% 6 64.8 1.3% 4 76.1 0.7% 5
Octo (Octo et al., 2024) 78.9 1.0% 5 85.7 0.9% 6 84.6 0.9% 5 51.1 1.3% 6 75.1 0.6% 6
$\pi_{0}$ (Black et al., 2024) 96.8 0.8% 2 98.8 0.9% 2 95.8 1.1% 2 85.2 1.2% 3 94.2 0.9% 2
$\pi_{0}$-FAST (Pertsch et al., 2025) 96.4 0.7% 3 96.8 0.7% 4 88.6 1.0% 4 60.2 1.4% 5 85.5 1.0% 4
GR00T N1 (Bjorck et al., 2025) 94.4 0.9% 4 97.6 1.0% 3 93.0 1.2% 3 90.6 1.0% 2 93.9 1.1% 3
@@
Table 3: Performance comparison with state-of-the-art policies on LIBERO Benchmark (Liu et al., 2023a).
@@
RT-1-X (OpenX-Embodiment et al., 2024) 0% 4.2% 0% 0% 1.1%
OpenVLA (Kim et al., 2024) 0% 0% 0% 4.1% 1.0%
SpatialVLA (Qu et al., 2025) 20.8% 20.8% 25.0% 70.8% 34.4%
Magma (Yang et al., 2025) 37.5% 29.2% 20.8% 91.7% 44.8%
$\pi_{0}$-FAST (Pertsch et al., 2025) 29.1% 21.9% 10.8% 66.6% 32.1%
Octo-Base (Octo et al., 2024) 12.5% 8.3% 0% 43.1% 16.0%
Octo-Small (Octo et al., 2024) 47.2% 9.7% 4.2% 56.9% 29.5%
RoboVLM (Li et al., 2024a) 45.8% 20.8% 4.2% 79.2% 37.5%
$\pi_{0}$ (Black et al., 2024) 83.8% 52.5% 52.5% 87.9% 69.2%
ThinkAct (Huang et al., 2025) 58.3% 37.5% 8.7% 70.8% 43.8%

### v5

# EO-1: An Open Unified Embodied Foundation Model for General Robot Control

@@ Abstract
The human ability to seamlessly perform multimodal reasoning and physical interaction in the open world is a core goal for general-purpose embodied intelligent systems. Recent vision-language-action (VLA) models, which are co-trained on large-scale robot and visual-text data, have demonstrated notable progress in general robot control. However, they still fail to achieve human-level flexibility in interleaved reasoning and interaction. In this work, we introduce EO-Robotics, consists of EO -1 model and EO -Data1.5M dataset. EO -1 is a unified embodied foundation model that achieves superior performance in multimodal embodied reasoning and robot control through interleaved vision-text-action pre-training. The development of EO -1 is based on two key pillars: (i) a unified architecture that processes multimodal inputs indiscriminately (image, text, video, and action), and (ii) a massive, high-quality multimodal embodied reasoning dataset, EO -Data1.5M, which contains over 1.5 million samples with emphasis on interleaved vision-text-action comprehension. EO -1 is trained through synergies between auto-regressive decoding and flow matching denoising on EO -Data1.5M, enabling seamless robot action generation and multimodal embodied reasoning. Extensive experiments demonstrate the effectiveness of interleaved vision-text-action learning for open-world understanding and generalization, validated through a variety of long-horizon, dexterous manipulation tasks across multiple embodiments. This paper details the architecture of EO -1, the data construction strategy of EO -Data1.5M, and the training methodology, offering valuable insights for developing advanced embodied foundation models.
@@ 1 Introduction
Generalist robot policies with open-world capabilities is essential for deploying autonomous robots in real scenarios, where they need to tackle diverse tasks, follow various instructions, adapt to different situations, and even handle unforeseen events. In the realm of digit domains, the paradigm of learning from large-scale multimodal datasets has demonstrated broad proficiency and generalization across various interactive and assistive applications. Extending this idea to the physical world, researchers are now training large vision-language-action models on extensive robotic datasets to realize generalist robots with similar versatility. Although recent efforts have demonstrated impressive progress in achieving dexterous manipulation and long-horizon physical control within constrained environments, open world generalization remains the central challenge for developing general-purpose autonomous AI agents in the physical world (intelligence2025pi_0_5, gemini_robotics). To address this challenge, robots must be capable of acquiring comprehensive world knowledge, engaging in human-level reasoning, and executing dexterous actions.
@@
Early generalist robot policies (kim2024openvla, black2024pi_0, pertsch2025fast) primarily extend vision–language models (VLMs) into vision-language-action (VLA) models with domain-specific robotic data, either through auto-regressive decoding of discrete action tokens or by incorporating additional continuous flow matching modules. However, because these VLA models are trained exclusively on robotic datasets, they are restricted to narrow task domains and specific environments. As a result, they suffer from diminished general semantic knowledge inherited from VLMs and exhibit limited instruction-following capabilities.
@@
Recently, several studies (intelligence2025pi_0_5, driess2025knowledge, lin2025onetwovla) have explored co-training VLA models with both web data and robotic data, showing promising generality when interacting with new objects and unseen backgrounds. Nevertheless, existing approaches remain predominantly generating robot action at the end of the VLA output sequence, overlooking the rich temporal dynamics and causal dependencies among vision, language, and action modalities inherent in open-world embodied interactions. Humans, in contrast, exhibit a flexible and interleaved synergy between multimodal embodied reasoning and physical action, enabling highly generalizable and dexterous manipulation in the open world, i.e., reasoning guides action and action results inform subsequent reasoning. This raises a fundamental research question: How can we design an effective training paradigm for generalist robot policies that support flexible and mutual-informed reasoning-acting integration?
@@
Inspired by the impressive results of advanced multimodal understanding and generation systems (deng2025emerging, ma2025janusflow, xie2025show, team2025nextstep), which demonstrate the superiority in interleaved multimodal pretraining, we propose a unified embodied model to enable flexible and powerful multimodal embodied reasoning and action generation through interleaved embodied pretraining. Achieving this requires not only an effective and unified architecture that excels in mix-modality generation, but also carefully structured multimodal embodied data that jointly integrate texts, images, videos, and robotic actions.
@@
Regarding model architecture, we adopt a single unified decoder-only transformer that integrates discrete auto-regressive decoding with continuous flow matching denoising. The unified model is built on a pre-trained VLM, thereby inheriting broad visual–language knowledge, and its shared parameters are further optimized with modality-specific objectives: next-token prediction for text and flow matching for robotic actions. The model applies causal attention cross the entire interleaved vision-text-action sequence to capture the sequential dependencies between reasoning and acting. In addition, two MLPs are added to encode and decode continuous robotic actions, complementing the original text and visual tokenizers. Unlike prior VLA models (black2024pi_0, intelligence2025pi_0_5, driess2025knowledge) that introduce extra action-specific modules to learn action generation, our design enables easier alignment between vision/language and action modalities through circumventing train new action-specific parameters from scratch, thereby achieving more effective cross-modal knowledge transfer in generalist robot policies.
@@ 2.1 Problem Formulation
where $V_{\pi_{\theta}}$ represents the parameters of $\pi_{\theta}$ to predict the vector field for robot action generation. The actions are generated by integrating the predicted vector field from $\tau=0$ to $\tau=1$, starting with random noise ${z}_{t}^{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and $\delta$ is the integration step size. The shared parameters in the unified transformer enable the seamless transfer of semantic knowledge from vision-language understanding to action generation, leading to more general reasoning capabilities and control capabilities in embodied scenarios (deng2025emerging, intelligence2025pi_0_5, driess2025knowledge).
@@ 2.2 Model and Training
The proposed unified model architecture processes interleaved multimodal inputs (i.e., text, images, videos, and actions) through a shared transformer backbone that generates both discrete token sequences and continuous action tokens. The model employs a text tokenizer and visual encoder to convert text and image patches into input tokens, while the robot state $\mathbf{q}_{t}$ is linearly projected into the same transformer embedding space, i.e., $x_{i}^{w},x_{i}^{I},x_{i}^{q}\in\mathbb{R}^{d}$. Note that the text tokenizer and visual encoder are inherited from pre-trained VLM and the state projector is random initialized. For flow matching action denoising, the input “noisy action” $\mathbf{a}_{t}^{\tau}$ is obtained by $\mathbf{a}_{t}^{\tau}=\tau\mathbf{a}_{t}+(1-\tau)z_{\tau}$, where $z_{\tau}\sim\mathcal{N}(0,I)$ is random noise. Then, another noisy action linear projector is utilized to embed noisy action and flow matching timestep $\tau$ into noisy action token $x_{i}^{a}\in\mathbb{R}^{d}$. The unified transformer backbone, initialized from Qwen 2.5 VL (qwen2_5_vl), processes this interleaved multimodal input sequence with shared parameters and generates output sequence through two separate heads: a language head for text token decoding and a flow head for continuous action denoising. Note that we do not introduce extra action-specific parameters to separately model action denoising as prior VLA models, enabling seamless integration of discrete multimodal embodied reasoning and continuous robot control.
@@
where $\mathbf{a}_{t}^{\tau}=\tau\mathbf{a}_{t}+(1-\tau)z_{0}$ is the input noisy action. $x_{<a}$ represents the all preceding multimodal tokens. As in (black2024pi_0), we sample the flow matching timestep $\tau$ from a beta distribution that emphasizes lower (noisier) timesteps. With the interleaved multimodal token sequence, we apply next token prediction $\mathcal{L}_{ar}(\theta)$ to the language head and flow matching $\mathcal{L}_{fm}(\theta)$ to the flow head for predicting denoising vector field. The unified model is trained end-to-end by optimizing sum of these two objectives: $\mathcal{L}=\mathcal{L}_{ar}(\theta)+\mathcal{L}_{fm}(\theta)$.
@@ 3 Dataset and Benchmark
LLaVA-1.5 (liu2024improved), LLaVA-Video-178K (zhang2024video),
PixMo-Points (molmo2024), RefCOCO (kazemzadeh-etal-2014-referitgame),
RoboVQA (sermanet2024robovqa)
@@
AgiBotWorld (bu2025agibot), Open X-Embodiment (o2024open),
RoboMIND (wu2024robomind), SO100-Community (shukor2025smolvla),
IPEC-Franka (qu2025spatialvla)
@@ 3.1 Data Source
Web Multimodal Data. The web text-image paired data is essential for multimodal understanding. We curated a web multimodal dataset comprising 5.7 Million samples with 7.1 Billion tokens, integrating six public sources across three task domains: i) Visual instruction following. We incorporate the LLaVA Visual Instruct Series (liu2023llava, sermanet2024robovqa) datasets, which contain GPT-generated multimodal instructions and robotics-oriented visual question answering examples. These sources cover diverse interaction patterns, including image-grounded conversation, multi-step reasoning, and long-horizon task planning. ii) Video understanding and reasoning. We leverage the LLaVA-Video (zhang2024llavanext-video) and RoboVQA (sermanet2024robovqa), which support temporal multimodal comprehension through detailed video captioning, open-ended QA, and multiple-choice question answering. iii) Referring expression and spatial grounding. To support fine-grained object understanding, we include RefCOCO (subset) (refcoco) for natural language referring expression comprehension and PixMo-Points (pixmo) for spatial localization via human-annotated coordinate grounding.
Robot Control Data. Robot control data provides critical supervision for policy learning across embodiment generalization and fine-grained manipulation. We aggregate a large-scale real robot control dataset comprising 1.2 Million episodes with 0.13 Trillion tokens from five public sources, forming one of the most comprehensive collections for robot learning to date. The dataset includes AgiBot-World (bu2025agibot), featuring dual-arm humanoid demonstrations across diverse bimanual manipulation tasks with rich visuomotor sequences; Open X-Embodiment (o2024open), which spans 22 robot types and 527 skills to support cross-platform generalization; and SO100-Community (shukor2025smolvla), a collection of long-horizon community demonstrations using SO100 platform. It also incorporates RoboMind (wu2024robomind), covering a wide range of tasks across multiple embodiments (e.g., Panda, UR5e, AgileX, humanoids) with diverse object classes.
@@ 4.1 Implementation Details
The EO -1 model is trained on a large-scale corpus that integrates 1.2M real-robot demonstrations from AgiBotWorld (bu2025agibot), Open X-Embodiment (o2024open), RoboMIND (wu2024robomind), and SO100-Community (shukor2025smolvla), together with 5.7M web multimodal samples and 1.5M interleaved embodied data pairs, yielding a total of 135B tokens. To optimize on this scale, we train for five epochs with Flash-Attention variable-length packing (average sequence length of 16,384) and a batch size of 1. The backbone learning rates are set to $5\times 10^{-5}$ for both the language model and multimodal projector, $1\times 10^{-5}$ for the vision encoder, and a fixed balance factor of $1.0$ is applied between language regression and action flow matching losses. Additional settings include an action denoising chunk size of $16$, resolution bounds of 50176–100352, and the DeepSpeed ZeRO-1 optimizer. At inference time, the policy is conditioned on multi-view camera observations and sub-task instructions. It predicts a 16-step action chunk through 10 denoising iterations for robot execution. This design achieves a balance between stability and efficiency, while requiring only 6 GB of GPU memory on a single NVIDIA RTX 4090, thus enabling real-time evaluation in both simulation and real-world environments.
@@ 4.2 Open-sourced Benchmark Evaluation Setups
  • RoboVQA (sermanet2024robovqa): A free-form visual question answering (VQA) benchmark designed to assess high-level reasoning in egocentric embodied scenes. Following prior works, model performance is evaluated using BLEU-4 scores.
  • ERQA (gemini_robotics): A curated benchmark that emphasizes spatial reasoning and grounded world knowledge. The tasks are designed around realistic robotic scenarios requiring inference over object relationships, geometry, and agent-centric spatial cues.
@@
  • SimplerEnv (li24simpler): A manipulation benchmark featuring WidowX and Google Robots operating in visually diverse environments with variations in lighting, surface textures, and camera viewpoints. It leverages the Bridge Data V2(walke2023bridgedata) for evaluating Real-to-Sim transfer. SimplerEnv includes both short-horizon atomic tasks and visually challenging variants, specifically designed to assess robustness under appearance shifts between real and simulated settings.
  • LIBERO (liu2023libero): A long-horizon manipulation benchmark consisting of temporally extended, multi-stage tasks across various object categories and interaction types. LIBERO emphasizes planning, memory, and semantic understanding in complex simulated environments.
@@ 4.3 Benchmark Evaluation Results
Embodied Reasoning. To evaluate the embodied reasoning capabilities, we benchmark it on three key tasks: RoboVQA sermanet2024robovqa, ERQA (gemini_robotics), and EO -Bench. These benchmarks assess a diverse range of reasoning skills, including spatial understanding, physical commonsense, and multistep task planning. We compare EO -1 against three categories of baselines: public VLMs (e.g., Qwen2.5 VL (qwen2_5_vl), InternVL2.5 (internvl)), private VLMs (e.g., GPT-4o, Gemini 1.5 Flash (gemini), Claude 3.5), and co-trained visual-language-action models (e.g., ChatVLA (zhou2025chatvla), Magma (yang2025magma)), with all models evaluated under consistent embodied conditions.
@@
Gemini 1.5 Flash (team2023gemini) 46.0 46.3 46.3 37.8 44.1
@@
Qwen2.5 VL 3B (qwen2_5_vl) 55.9 35.3 20.0 22.6 33.5
Qwen2.5 VL 7B (qwen2_5_vl) 56.9 39.3 26.7 31.5 38.6
InternVL2.5 4B (internvl2_5) 45.7 36.8 37.0 31.9 37.9
InternVL2.5 8B (internvl2_5) 43.9 45.2 32.8 38.1 40.0
ChatVLA 2B (zhou2025chatvla) 33.7 34.3 26.4 21.9 29.1
Magma 8B (yang2025magma) 30.3 29.3 29.4 36.7 31.4
@@
Table 2: Performance comparison with standard VLMs and co-trained visual-language-action models on RoboVQA (sermanet2024robovqa), ERQA (gemini_robotics), and EO -Bench benchmarks.
@@
OpenVLA (kim2024openvla) 84.7 0.9% 5 88.4 0.8% 5 79.2 1.0% 4 53.7 1.3% 5 76.5 0.6% 5
SpatialVLA (qu2025spatialvla) 88.2 0.5% 2 89.9 0.7% 4 78.6 0.6% 5 55.5 1.0% 4 78.1 0.7% 4
OpenVLA-OFT (kim2025fine) 97.6 0.9% 1 98.4 0.8% 1 97.9 1.0% 1 94.5 1.3% 1 97.1 0.6% 1
ThinkAct (huang2025thinkact) 88.3 –% 3 91.4 –% 3 87.1 –% 3 70.9 –% 3 84.4 –% 3
MolmoAct-7B-D (lee2025molmoact) 87.0 –% 4 95.4 –% 2 87.6 –% 2 77.2 –% 2 86.6 –% 2
Diffusion Policy (chi2024diffusionpolicy) 78.5 1.1% 6 87.5 0.7% 5 73.5 1.2% 6 64.8 1.3% 4 76.1 0.7% 5
Octo (team2024octo) 78.9 1.0% 5 85.7 0.9% 6 84.6 0.9% 5 51.1 1.3% 6 75.1 0.6% 6
$\pi_{0}$ (black2024pi_0) 96.8 0.8% 2 98.8 0.9% 2 95.8 1.1% 2 85.2 1.2% 3 94.2 0.9% 2
$\pi_{0}$-FAST (pertsch2025fast) 96.4 0.7% 3 96.8 0.7% 4 88.6 1.0% 4 60.2 1.4% 5 85.5 1.0% 4
GR00T N1 (bjorck2025gr00t) 94.4 0.9% 4 97.6 1.0% 3 93.0 1.2% 3 90.6 1.0% 2 93.9 1.1% 3
@@
Table 3: Performance comparison with state-of-the-art policies on LIBERO Benchmark (liu2023libero).
@@
RT-1-X (o2024open) 0% 4.2% 0% 0% 1.1%
OpenVLA (kim2024openvla) 0% 0% 0% 4.1% 1.0%
SpatialVLA (qu2025spatialvla) 20.8% 20.8% 25.0% 70.8% 34.4%
Magma (yang2025magma) 37.5% 29.2% 20.8% 91.7% 44.8%
$\pi_{0}$-FAST (pertsch2025fast) 29.1% 21.9% 10.8% 66.6% 32.1%
Octo-Base (team2024octo) 12.5% 8.3% 0% 43.1% 16.0%
Octo-Small (team2024octo) 47.2% 9.7% 4.2% 56.9% 29.5%
RoboVLM (li2023generalist) 45.8% 20.8% 4.2% 79.2% 37.5%
$\pi_{0}$ (black2024pi_0) 83.8% 52.5% 52.5% 87.9% 69.2%
ThinkAct (huang2025thinkact) 58.3% 37.5% 8.7% 70.8% 43.8%